OpenStudy (anonymous):

find the derivative of x^2 y^2+x^3+y=8 at (2,0)

myininaya (myininaya):

2xy^2+x^2(2yy')+3x^2+y'=0 can you solve this for y'

myininaya (myininaya):

y'(2yx^2+1)=-3x^2-2xy^2

myininaya (myininaya):

y'=[-3x^2-2xy^2]/[2yx^2+1]

OpenStudy (anonymous):

thanks so now do I just plug in 2 for x and 0 for y?

myininaya (myininaya):

yes

OpenStudy (anonymous):

The derivative \[2xy^2 + 2x^2y(y') + 3x^2 + y' = 0\] So rearrange giving: \[y'(1+ 2x^2y) = -2xy^2 - 3x^2\] as myinaya said

OpenStudy (anonymous):

thanks guys!